### v1

# thanks: These authors co-supervised the projectthanks: These authors co-supervised the project

Abstract
Dissipation engineering offers a powerful tool for quantum technologies. Recently, new superconducting devices demonstrated an engineered two-photon dissipation rate exceeding all other relevant timescales. In particular, they have proven most useful to prevent transitions between the logical states |±$|\pm\alpha\rangle$ of a cat qubit. Here, we present three key applications of strong two-photon dissipation for quantum measurement and control, beyond cat qubit stabilization. Firstly, we demonstrate its efficacy in overcoming limitations encountered in Wigner tomography at high photon numbers. Secondly, we showcase its potential for realizing universal gates on cat qubits, exploiting the coherent mapping between cat qubit states and superpositions of 0 and 1 photons. Finally, we harness the transient dynamics of a cat state under two-photon dissipation to prepare squeezed cat states with a squeezing factor exceeding 3.8 dB.
The perception of dissipation in quantum systems has undergone a drastic change. Once seen as a relentless foe, it is now being recognized as a potential ally, capable of being ‘tamed’ to become a valuable tool for quantum measurements, control, and quantum error correction [1, 2, 3, 4, 5]. Within the framework of bosonic codes, dissipation offers a strategy to remove entropy from a memory mode (e.g., superconducting resonator, trapped ions, spin ensemble) and thereby stabilize a logical qubit [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. This is achieved by carefully tailoring the coupling between the memory and its environment, inducing an effective dissipation that confines information within the desired qubit subspace and prevents leakage into the broader Hilbert space. This technique has proven particularly successful in stabilizing cat qubits[8, 9, 11, 12, 16, 18, 19], whose computational basis comprises two coherent states |±$\ket{\pm\alpha}$. A key resource in this process is two-photon dissipation, where the memory loses pairs of photons to a low-temperature environment at a rate $\kappa_{2}$. A two-photon drive acting on the memory completes the stabilization process of the cat qubit.
This work explores the broader potential of two-photon dissipation beyond cat qubit stabilization. We leverage the autoparametric cat device [19], which boasts an engineered two-photon dissipation rate $\kappa_{2}$ exceeding the residual single-photon loss rate $\kappa_{1}=(11~{}\mu\mathrm{s})^{-1}$ by two orders of magnitude. Following proposals in [22, 23], we harness two-photon dissipation to extend Wigner tomography protocols to large photon numbers, and achieve universal control of cat qubits within a timescale $\kappa_{2}^{-1}$. Furthermore, we demonstrate the preparation of squeezed cat states with a squeezing parameter reaching 3.9 dB.
Existing devices [9, 11, 12, 16, 18, 19] that demonstrate two-photon dissipation use a dedicated mode, denoted as the buffer, to mediate the interaction between the memory and its environment. In our experiment, this buffer dissipates photons at a rate $\kappa_{b}/2\pi\approx 40~{}\mathrm{MHz}$ (Fig. 1a). The memory and buffer modes are coupled through a nonlinear element implementing a two-to-one photon exchange Hamiltonian $\hat{H}=\hbar g_{2}\hat{m}^{2}\hat{b}^{\dagger}+h.c.$ at a rate $g_{2}$. Here, $\hat{m}$ and $\hat{b}$ represent the annihilation operators of the memory and buffer modes, respectively. This interaction is realized in our experiment using a 3-wave mixing interaction under the condition $2\omega_{m}=\omega_{b}$, characteristic of autoparametric systems [19, 24]
Operating in the regime $\kappa_{b}\geq 8g_{2}|\alpha|$ ensures that photons in the buffer mode are evacuated into the environment at a faster rate than they are converted into pairs of photons in the memory. This allows for the adiabatic elimination of the buffer mode, leading to an effective dissipation operator $\sqrt{\kappa_{2}}\hat{m}^{2}$ where $\kappa_{2}=4g_{2}^{2}/\kappa_{b}$. In the autoparametric cat device [19], an external flux bias $\phi_{\mathrm{ext}}$ can be used to activate or deactivate two-photon dissipation. We have identified a specific flux point, $\phi_{\mathrm{on}}$, at which $\omega_{b}=2\omega_{m}$, resulting in $\kappa_{2}/2\pi=2.16~{}\mathrm{MHz}$. Another flux point, $\phi_{\mathrm{off}}$ with negligible $\kappa_{2}/2\pi$ serves to disable two-photon dissipation (details in [24]). It is chosen at a point where the self-Kerr effect is weak.
Figure 1: (a) Schematic of a typical device for two-photon dissipation. The memory mode (blue) interacts with the buffer mode (green) via a 2-to-1 photon exchange with coupling strength $g_{2}$. The buffer can be driven with an amplitude $\epsilon_{d}$. Additionally, a dispersively coupled tomography transmon (orange) with frequency $\omega_{q}$ can be read out through a dedicated readout resonator at $\omega_{r}$. (b) Photon number distribution $\mathbb{P}_{n}$ of an arbitrary quantum state before and after undergoing two-photon dissipation.
Let us begin by demonstrating how two-photon dissipation can address limitations in Wigner tomography, as proposed in [23]. The Wigner function of a density matrix $\hat{\rho}$ is proportional to the expectation value of the parity operator $\hat{\mathcal{P}}=e^{i\pi\hat{m}^{\dagger}\hat{m}}$after the memory has undergone a displacement operation $\hat{\mathcal{D}}\left(-\beta\right)=e^{-(\beta\hat{m}^{\dagger}-\beta^{*}\hat{m})}$:
$W_{\hat{\rho}}\left(\beta\right)=\frac{2}{\pi}\mathrm{Tr}\left(\hat{\mathcal{D}}\left(\beta\right)^{\dagger}\hat{\rho}\,\hat{\mathcal{D}}\left(\beta\right)\mathcal{\hat{P}}\right).$ (1)
The displacement is achieved by driving the memory on resonance with a pulse whose complex amplitude scales with $-\beta$. In our system, the pulse duration of $20~{}\mathrm{ns}$ is chosen to be significantly shorter than the inverse self-Kerr rate of the memory $\chi_{\mathrm{m,m}}\approx(720~{}\mathrm{ns})^{-1}$. Wigner tomography then boils down to measuring the average value of $\hat{\mathcal{P}}$.
Various parity measurement methods have been demonstrated in superconducting circuits, usually relying on coupling the memory at resonance [25, 26] or dispersively[27, 28, 29, 30, 31] to an auxiliary qubit. However, all these methods deteriorate as the average photon number $\bar{n}$ after displacement increases. Let us focus on the Ramsey-like sequence proposed in [27]. The qubit is first prepared in an equal superposition of ground and excited state $(|g\rangle+|e\rangle)/\sqrt{2}$ by applying an unconditional $\pi/2$ pulse, much shorter than the inverse dispersive coupling, which is $(\chi_{\mathrm{q,m}})^{-1}\approx 0.9~{}\mu\mathrm{s}$ in our experiment. After a time $T_{\mathrm{parity}}=\pi/\chi_{\mathrm{q,m}}$, the superposition accumulates a phase $n\pi$ with $n$ the number of photons of the memory. Even and odd numbers of photons are then respectively mapped onto the states (|$\left(\ket{g}+\ket{e}\right)/\sqrt{2}$ and (|$\left(\ket{g}-\ket{e}\right)/\sqrt{2}$ of the qubit. A final unconditional $\pi/2$ pulse is applied, mapping the memory parity operator $\hat{\mathcal{P}}$ onto the Pauli operator $\hat{\sigma}_{z}=|e\rangle\langle e|-|g\rangle\langle g|$ of the qubit, which can then be readout.
Firstly, this protocol underestimates the amplitude of the Wigner function since the parity outcome is flipped when a photon is lost during the idle time $T_{\mathrm{parity}}$, which occurs with probability $1-e^{-\kappa_{1}T_{\mathrm{parity}}\bar{n}}$. Additionally, for large $\bar{n}$, the superconducting qubit state can leak out of its subspace $\mathrm{span}(|g\rangle,|e\rangle)$, especially after the qubit gets excited with the $\pi/2$-pulse [32, 33, 34, 35, 36, 37]. Finally, the upper bound on the $\pi/2$ pulse duration scales as $(\chi_{\mathrm{q,m}}\bar{n})^{-1}$, requiring impractically high driving power at large photon numbers. These limitations at high photon numbers can be overcome using the protocol proposed in Ref. [23]. The key lies in leveraging two-photon dissipation’s ability to preserve parity while transforming a state with an arbitrary photon number distribution into a state with at most $1$ photon (Fig. 1b).
To illustrate typical errors of Wigner tomography, we prepare a cat state |$\ket{C_{\alpha}^{+}}\propto|\alpha\rangle+|-\alpha\rangle$ and realize its tomography. The preparation is performed by tuning the flux to $\phi_{\mathrm{on}}$, which turns on the two-photon dissipation, and driving the buffer mode on resonance for 300 ns with a drive amplitude $\epsilon_{d}=-g_{2}^{*}\alpha^{2}$. As demonstrated in Refs. [9, 19], this implements an effective dissipation operator $\hat{L}_{2}=\sqrt{\kappa_{2}}\left(\hat{m}^{2}-\alpha^{2}\right)$. Fig. 2b shows the obtained Wigner tomography for $\alpha=3.3$. The typical fringes of cat states can be seen close to $\beta=0$ but their contrast is much smaller than expected for |$\ket{C_{\alpha}^{+}}$. Part of this loss of contrast comes from single photon losses during preparation. However, in the cut $W_{\hat{\rho}}(\beta)$ along $\beta\in i\mathbb{R}$ (Fig. 2.d), one can see that the measured Wigner function (orange dots) does not faithfully reproduce numerical simulations only accounting for the preparation infidelity (grey line). Here, most of the tomography errors can be attributed to single photon loss during parity measurement. Indeed, the probability of losing a photon during $T_{\mathrm{parity}}\approx 2.7~{}\mu\mathrm{s}$ reads $1-e^{-0.26\bar{n}}$ when the memory has $\bar{n}$ photons on average, reaching $50~{}\%$ for 2.7 photons. The error is therefore dramatic for the cat state in Fig. 2.b which has more than 10 photons on average. This effect is reproduced by simulations (orange line in Fig. 2.d). In contrast, the measurement of $W_{\hat{\rho}}(\pm\alpha)$ is only marginally affected by single photon loss. Indeed, after a displacement $\hat{\mathcal{D}}\left(\pm\alpha\right)$, the memory ends up in a superposition of $|0\rangle$ and |±$|\pm 2\alpha\rangle$. The parity of $|0\rangle$ does not depart from 1 during measurement and the parity of |±$|\pm 2\alpha\rangle$ is negligible, leading to a measured ${\hat{\rho}}(\pm\alpha)\approx 1/\pi$ as expected.
We solve this issue by quickly turning the two-photon dissipation back on by tuning the flux to $\phi_{\mathrm{on}}$ after the displacement (Fig. 2a). Idling at $\phi_{\mathrm{on}}$ for $300~{}\mathrm{ns}$ (a few $\kappa_{2}^{-1}$) then ensures that the memory decays to a state with at most 1 photon while preserving its parity (Fig. 1b). Tuning the flux back to $\phi_{\mathrm{off}}$ enables parity measurement. In our example of a cat state above, a better measurement of the Wigner fringes can then be achieved (Fig. 2c,d). Simulations corresponding to this parity measurement protocol (blue line) predict a drastic improvement of the fringe reconstruction, reproducing what is expected when only accounting for preparation infidelity (grey line). This is confirmed by the measured Wigner function (blue dots) which shows an increase in contrast compared to the usual parity measurement protocol. It should be noted that avoiding single-photon loss while removing pairs of photons requires a large $\kappa_{2}/\kappa_{1}$ ratio. This condition can be met with other devices than the autoparametric cat, such as a simple parametrically pumped transmon qubit [11], making this method accessible to a wide range of experiments using superconducting qubits.
Figure 2: (a) Pulse sequence of the dissipation enhanced Wigner tomography of |$\ket{C_{\alpha}^{+}}$ including an extra 300 ns idle time at $\phi_{\mathrm{ON}}$ after memory displacement. (b) Left: Wigner tomography of |$\ket{C_{\alpha}^{+}}$, $\alpha=3.3$, without the extra idle time. (c) Wigner tomography of |$\ket{C_{\alpha}^{+}}$ using the dissipation enhancement. (d) Cuts of the Wigner function along the imaginary axis $\mathrm{Re(\beta)}=0$ in various scenarios. Solid gray line: simulated memory state at the end of the 300 ns buffer pulse $\epsilon_{d}$. Orange points: Wigner tomography in (b). Solid orange line: simulation of the measurement protocol without the extra idle time. Blue points: Wigner tomography in (c). Solid blue line: simulation of the dissipation enhanced measurement protocol.
Two-photon dissipation can also be used to control cat qubits. For instance, driving the memory on resonance while the dissipation $\hat{L}_{2}$ is turned on performs a gate $\hat{Z}(\theta)$ [11, 19, 18] . In contrast, $\hat{X}(\theta)$ and $\hat{Y}(\theta)$ gates are seemingly incompatible with the engineered dissipation $\hat{L}_{2}$, which prevents population transfer between $|\alpha\rangle$ and $|-\alpha\rangle$. However, one can use the coherent mapping between the manifolds $\mathrm{span}\left(\ket{0},\ket{1}\right)$ and $\mathrm{span}\left(\ket{\alpha},\ket{-\alpha}\right)$ to perform these gates while the memory transits through the $\{|0\rangle,|1\rangle\}$ manifold (Fig. 3a) [22]. These deflation and inflation processes are performed by adiabatically modulating the buffer drive between $|\epsilon_{d}|=0$ and $|\epsilon_{d}|=g_{2}\left|\alpha\right|^{2}$, so that the memory follows a steady state of the dissipation. From this point of view, the preparation of |$\ket{C_{\alpha}^{+}}$ in Fig. 2 corresponds to the inflation of the vacuum state, and the Wigner tomography is enhanced by deflation before parity measurement.
Figure 3: (a) Inflation and deflation processes coherently map the manifolds $\mathrm{span}\left(\ket{\alpha},\ket{-\alpha}\right)$ and $\mathrm{span}\left(\ket{0},\ket{1}\right)$ by turning on or off the buffer drive $\epsilon_{d}$ [22, 18]. (b) Pulse sequence used to prepare an initial state |$\ket{\alpha}$, apply the holonomic $\hat{X}\left(\theta\right)$ gate, and measure the memory Wigner function. (c) Schematic evolution in the memory phase space of a coherent state while applying the holonomic $\hat{X}\left(\theta\right)$ gate. (d) Measured evolution of the memory Wigner function, starting from the coherent state |$\ket{-2.27}$, during the $\hat{X}\left(\pi/2\right)$ gate. The last snapshot occurs right before the final virtual rotation $\hat{R}(-\pi/2)$. An animation showing every measured time can be found as supplementary material [24].
We implement a holonomic $\hat{X}\left(\theta\right)$ gate by modulating the buffer drive amplitude $\epsilon_{d}(\alpha)$ while remaining at $\phi_{\mathrm{on}}$ (Fig. 3b). We first turn off the buffer drive to deflate the memory state into the $\{|0\rangle,|1\rangle\}$ manifold. The buffer drive is then turned back on with an additional phase $2\theta$, $\epsilon_{d}\rightarrow\epsilon_{d}\left(\alpha\right)e^{2i\theta}$ [24], mapping the memory into the $\{|\alpha e^{i\theta}\rangle,|-\alpha e^{i\theta}\rangle\}$ manifold. This additional phase implements a virtual $\hat{X}(\theta)$ gate in the $\{|0\rangle,|1\rangle\}$ manifold (see Fig. 3a). A counter-rotation $\hat{R}\left(-\theta\right)$ is finally applied to bring the memory back into the cat code while preserving the phase of the superposition. This final rotation could be performed by adiabatically sweeping the phase of the buffer drive back to 00. Equivalently, we choose to apply an instantaneous virtual operation by changing the phase of the local oscillator of the memory drive by $-\theta$. The evolution of the memory in phase space starting from |$\ket{-\alpha}$ is represented in Fig. 3c. We probe the evolution of the memory during the $\hat{X}\left(\pi/2\right)$ gate starting from the coherent state |$\ket{-\alpha}$ with $\alpha=2.27$. This particular gate has been used for tomography without a transmon in [18]. Wigner tomographies of the memory are shown in Fig. 3d at $6$ different times. The memory state can first be seen evolving from the coherent state towards (|$\left(\ket{0}-\ket{1}\right)/\sqrt{2}$ (third snapshot) before being mapped into (|$\left(\ket{\alpha e^{i\pi/2}}-i\ket{-\alpha e^{i\pi/2}}\right)/\sqrt{2}$ (last snapshot).
We show in Figs. 5 of [24] the state of the memory after applying the $\hat{X}\left(\theta\right)$ gate on the same coherent state |$\ket{-\alpha}$ for an arbitrary angle $\theta$. Note that the shape of the buffer drive during the gate was optimized with numerical simulations shown in Fig. S3 of [24]. The gate error being substantial, we settle for characterizing the gate properties by its effect on the coherent state $|-\alpha\rangle$ only. We compute the trace distance $T\left(\hat{\rho},|\psi(\theta)\rangle\langle\psi(\theta)|\right)$ [38, 24] between the density matrix $\hat{\rho}$ reconstructed from the measured Wigner function and the state $|\psi(\theta)\rangle=\hat{X}\left(\theta\right)|-\alpha\rangle$. The largest distance occurs at $\theta=\pm\pi/2$, where it reaches $0.23$ (see Fig. S4 in [24]). We attribute part of this apparent error to the remaining measurement infidelity (deviations between blue dots and grey line in Fig. 2d). For $\theta$ close to multiples of $\pi$, the Wigner tomography of a coherent state does not lead to such measurement errors and simulations of the memory evolution during the gate reproduce the observed trace distance. Additionally, simulations with and without the dephasing rate $\kappa_{\varphi}^{\mathrm{m}}/2\pi=0.08~{}$MHz reveal it has a major contribution to the gate error owing to the induced decoherence when the memory is in the $\{|0\rangle,|1\rangle\}$ manifold. The remaining gate errors can be attributed to a non-adiabatic evolution of the memory during the drive, or single photon loss.
Figure 4: Preparation of a squeezed cat state. (a) Wigner functions of the memory at early times (0, 16 ns, or 24 ns) during the deflation of |$\ket{C_{\alpha}^{+}}$, where $\alpha=2.9$. The measurements (left) are compared with numerical simulations made without free parameters (right). (b) Schematic representation of the buffer and memory states, in the non-adiabatic regime, at the beginning of deflation. An effective displacement of mode $\hat{b}$ creates an effective squeezing Hamiltonian acting on the memory during its decay to the {|$\left\{\ket{0},\ket{1}\right\}$ manifold.
Interestingly, probing the deflation dynamics at short times (a few $\kappa_{b}^{-1}$) reveals a potential resource for state preparation. The deflation spontaneously prepares squeezed cat states. After preparing the state |$\ket{C_{\alpha}^{+}}$ in the memory, the drive $\epsilon_{d}(\alpha)$ is turned off while the flux remains at $\phi_{\mathrm{on}}$. The measured Wigner functions of the memory state are shown in Fig. 4a for three different times between 0 and 24 ns. As the memory releases photon pairs to the buffer, it acts as an effective drive on the buffer with an amplitude $g_{2}\hat{m}^{2}\right>=g_{2}\alpha^{2}$. This drive, which should not have an impact on the buffer when $\kappa_{b}\gg 8g_{2}\alpha$, here displaces the buffer into a state whose mean amplitude $\langle\hat{b}\rangle$ is proportional to $\alpha^{2}$. In turn an effective squeezing Hamiltonian
$\hat{H}_{\mathrm{m,eff}}/\hbar=g_{2}\langle\hat{b}\rangle\hat{m}^{\dagger 2}+h.c$ (2)
acts on the memory, which compresses the superposition of coherent states (Fig. 4b). As the memory deflates, it reaches a maximal squeezing at a time that decreases with the initial photon number. This observation is compatible with the expected evolution of the squeezing parameter $r$ at small time in $(g_{2}|\alpha|t)^{2}$ [24]. From there, the buffer decays to vacuum, hence suppressing the effective squeezing Hamiltonian. To extract the squeezing ratio $e^{2r}$, we fit the two measured peaks in the Wigner function with Gaussian functions of identical variance. We observe squeezing ratios as high as $e^{2r}\approx 3.9~{}\mathrm{dB}$ [39], obtained after a time $t=16$ ns, starting from |$\ket{C_{\alpha}^{+}}$ with $\alpha=2.9$ (Fig. 4a). Note that during the evolution of the memory, the product of the largest and smallest variances remains close to that of the vacuum state. The predicted evolution using the bipartite buffer-memory system, without the adiabatic elimination, reproduces qualitatively the measured Wigner functions (right panels). A rounding of the Wigner function appears in the theory owing to two-photon dissipation rate, which ultimately overcomes the squeezing rate as the buffer is depleted. While bigger compression is expected starting with larger values of $\left|\alpha\right|$, an uncontrolled deformation of the initial cat state in the experiment limits the squeezing factor. We attribute this deformation and the discrepancy between simulations and experiments to the self-Kerr, the dephasing of the buffer or a detuning between $2\omega_{m}$ and $\omega_{b}$.
This work illustrates how two-photon dissipation is a remarkable tool for quantum measurement and control. Let us put in perspective the three applications we demonstrated. The enhancement of Wigner tomography is straightforward for any device with $\kappa_{2}\gg\kappa_{1}$ and should be useful for a broad range of experiments. Furthermore, the demonstration of an arbitrary $\hat{X}(\theta)$ gate together with the known $\hat{Z}(\theta)$ gate [11, 19, 18] provides universal control of cat qubits. We note that since $\hat{X}(\pi)$ gate can be performed virtually by changing the phase of the local oscillator of the memory, such a control is not required for quantum error correcting codes [40]. However, our experiment in Fig. 3 provides a useful tool to explore and optimize the dynamics of the holonomic gate, which is an essential component of cat qubit readout without transmon qubits [18]. Finally, two-photon dissipation appears to be an interesting approach for the preparation of squeezed cat states, whose coherence properties are promising for quantum information processing [41, 42, 43, 17, 44].
Acknowledgements.
This research was supported by the QuantERA grant QuCos ANR-19-QUAN-0006, the Plan France 2030 through the project ANR-22-PETQ-0003. We acknowledge IARPA and Lincoln Labs for providing a Josephson Traveling-Wave Parametric Amplifier. We thank Zaki Leghtas and Mazyar Mirrahimi for inspiring discussions and feedback.